OpenStudy (anonymous):

(x + 1)2 + (y – 2)2 = 16 I'm suppose to find the radius and center of each circle, can someone please show me how?

OpenStudy (anonymous):

the general form is \[(x-h)^2+(y-k)^2=r^2\] for a circle with center at \((h,k)\) and radius \(r\)

OpenStudy (anonymous):

ok so how do i find it?

OpenStudy (anonymous):

match up the general to the specific to get your answer \[(x-h)^2+(y-k)^2=r^2\\ (x+1)^2+(y-2)^2=4^2\]

OpenStudy (anonymous):

your job is to identify \(h\), \(k\) and \(r\)

OpenStudy (anonymous):

@satellite73 how?

OpenStudy (anonymous):

pretty much with your eyeballs. \[(x-\color{red}h)^2\\(x-\color{red}{(-1)})^2\] tells you \(\color{red} h =\color{red}{ -1}\)

OpenStudy (anonymous):

\[(y-\color{blue}k)^2\\(y-\color{blue}2)^2\] what is \(\color{blue}k\) ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Luigi0210 a little help please?

OpenStudy (luigi0210):

Use satellite's formula

OpenStudy (anonymous):

yes, you are right, \(k=2\)

OpenStudy (anonymous):

how do i find the radius?

OpenStudy (anonymous):

and for the radius \[16=r^2\\ \color{green}4^2=\color{green}r^2\]\]

OpenStudy (anonymous):

so \(\color{green}r=\color{green}?\)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!!!!!